ROVIBRONIC BANDS OF THE Ã ← X̃ TRANSITION OF CH_{3}OO and CD_{3}OO DETECTED WITH CAVITY RINGDOWN NEAR 1.2 μM

\underline{CHAO-YU CHUNG}, Department of Chemistry, National Tsing Hua University, Hsinchu 30013, Taiwan; CHI-WEN CHENG and YUAN-PERN LEE, Department of Applied Chemistry and Institute of Molecular Science, National Chiao Tung University, Hsinchu 30010, Taiwan and Institute of Atomic and Molecular Sciences, Academia Sinica, Taipei 10617, Taiwan.

We have recorded several rovibronic bands of the Ã ← X̃ transition of CH_{3}OO and CD_{3}OO near 1.2 μm with the cavity ringdown technique. The light source was a Raman-shifted beam from a dye laser pumped with a frequency-doubled Nd:YAG laser. The 7_{0}^{1} band shows partially resolved rovibronic lines which, when coupled with rotational parameters of the lower state determined recently by Endo and coworkers, can be fitted to yield spectral parameters of the upper states: ν = 8269.6 ± 0.5 cm^{−1}, A′ = 1.526 cm^{−1}, B′ = 0.388 cm^{−1}, and C′ = 0.323 cm^{−1} for CH_{3}OO, and ν = 8195.5 ± 0.5 cm^{−1}, A′ = 1.169 cm^{−1}, B′ = 0.329 cm^{−1}, and C′ = 0.28 cm^{−1} for CD_{3}OO. The vibrational spacing of 886.8 cm^{−1} for the ν_{7} mode of CH_{3}OO is more precise than a previous report of 896 ± 9 cm^{−1} using conventional absorption spectroscopy; this mode is associated with the C−O stretching mode rather than the O−O stretching mode assigned previously. Two additional vibronic bands of CD_{3}OO were also detected for the first time; tentatively identified vibrational wave numbers of the Ã state of CD_{3}OO are 953 (ν_{6}) and 970 (ν_{5}) cm^{−1}, consistent with quantum-chemical calculations. Hot bands associated with excitation of the torsional mode were also observed.